OpenStudy (anonymous):

Which of the following is not an element of the following set: A: {positive integers divisible by 4} A) 8 B) 18 C) 48 D) 12

OpenStudy (anonymous):

8/4 = 2 so 8 cannot be the answer

OpenStudy (anonymous):

12/4 = 3 so 12 cannot be the right answer

OpenStudy (anonymous):

you try the other two, and the one that you cant evenly divide by four is the answer

OpenStudy (anonymous):

@tjones89 so would 48 be the correct answer?

OpenStudy (anonymous):

48 is not the right answer

OpenStudy (anonymous):

@tjones89 so 18 is the right answer?

OpenStudy (anonymous):

lets find out. If a number is evenly divisible by four, I should be able to add 4 to itself until I get that number for example 4 + 4 + 4 = 12 so 12 is evenly divisible by 4

OpenStudy (anonymous):

4 + 4 + 4 + 4 = 16 so 16 is divisible by 4

OpenStudy (anonymous):

4 + 4 + 4 + 4 + 4 = 20 so 20 is divisible by 4, but what happened to 18? We skipped over it, so it must not be divisible by 4

OpenStudy (anonymous):

so it would have to be 12, because its evenly divisible by four

OpenStudy (anonymous):

nope, the question is asking for a number that is NOT divisible by four

OpenStudy (anonymous):

which of the following is NOT an element of the set...

OpenStudy (anonymous):

so its 18?

OpenStudy (anonymous):

@tjones89 so its 18?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@tjones89 thank you! :)